### v1

# ADVANCING MATHEMATICS RESEARCH WITH LARGE LANGUAGE MODELS

Abstract.
The main drawback of using generative AI for advanced mathematics via Large Language Models (LLMs) is that they are probabilistic pattern-matchers, not logical reasoning engines. However, LLMs can pick up on patterns in higher mathematics that are difficult for humans to see. By putting the design of LLMs to their advantage, mathematicians may use them as powerful interactive assistants that can carry out laborious tasks, generate and debug code, check examples, formulate conjectures and more. We discuss how LLMs can be used to advance mathematics research by careful use of prompt engineering. We also discuss the integration of LLMs with Computer Algebra Systems and formal proof assistants such as Lean.
1. Introduction
Mathematicians have mixed views about the role of LLMs in the mathematical landscape. While these models can look convincingly like they replicate known mathematics, on careful scrutiny, they fall short of the standards of rigor that the field requires.
This shortfall is built into the design of LLMs. They are fundamentally statistical, not logical, engines. As next-word predictors, they lack a built-in engine for formal symbolic deduction, meaning their emergent ‘reasoning’, called Natural Language Reasoning, is derived from statistical patterns in their training data.
However, this statistical foundation also provides certain benefits, as LLMs can pick up on patterns in higher mathematics that are difficult for humans to see. In particular, they have a learned geometric representation of mathematical language. By leveraging these capabilities and working with LLMs as they were designed to operate, mathematicians can use them as powerful interactive assistants.
In Section 2, we discuss the mechanics of probabilistic language generation. In Sections  37, we explore the foundational capabilities and limitations of LLMs, the nature of their training data, and their ability to detect geometric patterns in mathematical language. In Sections 815, we discuss applications and user techniques. This includes examples from combinatorial group theory, a comparison of current models, and a guide to prompt engineering, fact-checking and managing the context window. In Section 16 we discuss the integration of LLMs with Computer Algebra Systems and formal proof assistants like Lean. In Section 17, we give some examples of advancements in research in infinite dimensional group theory obtained with the strategic use of hybrid methods.
2. The mechanics of probabilistic language generation
2.1. The LLM as a probabilistic knowledge graph
Consider the following simplified view of an LLM as a ‘knowledge graph’ which represents all possible connections and probabilities within the LLM’s knowledge base.
Each word is a node in the graph. Each edge is a statistical relationship or an association between words, which cluster to form concepts.
additionnegative numbersirrational numbersintegersreal numbersinverseCONVERGED PROBABILITY
Weighted edges between trillions of nodes represent how words or concepts are related. The weights are probabilities indicating the likelihood of connection.
An LLM generates text by predicting the next word, while calculating the probabilities of all possible next nodes. It chooses one of the most likely paths through the knowledge graph.
2.2. Tokens: the building blocks of language
In AI, a token is the fundamental building block of language. AI models convert input materials (prompts) into tokens. A token $t_{i}$ is a digital representation of a word, a symbol, punctuation, a diagram, or other input.
Let $V$ be a finite vocabulary consisting of token representations of the training data for an LLM. An input to an LLM is a sequence of tokens $(t_{1},t_{2},\dots,t_{n})$. An output of an LLM is a sequence of tokens $(t_{1},t_{2},\dots,t_{m})$, where each $t_{i}\in V$.
The fundamental operation of an LLM is to compute the probability of the next token, $t_{k+1}$, conditioned on the sequence of all preceding tokens $(t_{1},t_{2},\dots,t_{k})$.
All AI language models have a maximum token limit for input and output.
Model Maximum Input Limit (tokens) Maximum Output Limit (tokens)
Gemini 1,048,576 65,536
ChatGPT 272,000 128,000
Claude 200,000 32,000
The token input limit has an impact on the work possible in a single chat session.
2.3. How the prompt shapes an LLM’s output
The prompt provides an initial sequence of tokens that the model uses to condition its output. The LLM generates text by sequentially examining the existing sequence of tokens, assigning a probability to every possible next token, and then choosing one to add to the sequence.
The probability of each potential next token is calculated based on the full sequence of tokens that came before it – both the original prompt and any text already generated.
The final output is the result of a chain of these probabilistic decisions. The framing of the initial prompt is critical in determining the usefulness of the output.
The design of prompts, or ‘prompt engineering’, significantly influences the behavior of an LLM [PromptGuide], [PromptEngineering]. It disambiguates the input, leading to more relevant and useful output.
3. Drawbacks of using generative AI in mathematics
The primary drawback of using current LLMs for mathematics is that they are probabilistic pattern-matchers, not logical reasoning engines. Their core function is to predict the next most likely word or symbol in a sequence, not to apply deterministic mathematical rules. An LLM is an expert at reconstructing what a proof should ‘look like’, based on every proof it has ever seen. This limitation leads to several key problems:
Mathematical hallucinations: LLMs generate outputs that look plausible and are often formatted correctly but contain nonsensical logic, invented theorems, or critical errors. They can state these falsehoods with the same confident authority as factual information.
Lack of true reasoning: Unlike a Computer Algebra System, an LLM cannot manipulate symbolic expressions according to deterministic rules. This makes it prone to subtle but serious errors in algebra, analysis, arithmetic, and logic. It is fundamentally incapable of handling verification.
Propagation of errors: LLMs learn from their training data, which includes errors and misconceptions found online. For example, it is known that ChatGPT 4.0 contains training data from retracted scientific papers [MITTR2025]. An LLM will inherit and reproduce these mistakes, making them unreliable for tasks that require rigor.
Consequently, mathematical output from an LLM has to be carefully scrutinized for logical and symbolic correctness.
4. Mathematical training data for LLMs
Training data for LLMs includes web documents, code with mathematical content, textbooks, online faculty-authored lecture notes and course materials, solutions to problem sets, as well as academic and scientific journal papers and content from arXiv. Google’s data collection from Google Books, for example, is unparalleled in its scale.
The training data includes the standard undergraduate mathematics curriculum from US and other universities, as well as the standard coursework curriculum for PhD programs in mathematics from US and other universities.
Differences in the content and handling of training data of AI companies are primary reasons why various AI models have distinct strengths, weaknesses, and ‘personalities’, especially in a specialized domain like mathematics. Curation and filtering of data, such as removing low-quality content and emphasizing trusted sources, is essential. However, the training material for each model and its handling remains a closely guarded trade secret.
A key limitation is that an LLM’s knowledge base is heavily biased towards materials that are easily digitized.
5. LLMs simulate computer algebra systems
LLMs simulate a scientific computing environment by delegating the computation to specialized Python libraries.
Table 1. Key Python libraries used by LLMs for mathematics
Library Main purpose Key capabilities & LLM use
NumPy Numerical Computing Handles arrays & matrices. Used by LLMs for linear algebra (matrix multiplication, row reduction, eigenvalues).
SymPy Symbolic Mathematics Performs precise algebra & calculus. Used for symbolic derivatives, integrals, and simplifying expressions.
SciPy Scientific Computing Provides advanced numerical routines. Used for optimization (finding minima/maxima) and solving differential equations.
Pandas Data Analysis Manages structured, table-like data. Used for reading and analyzing data from files such as Excel.
Matplotlib Plotting & Visualization Creates a wide variety of 2D graphs and charts. Used to plot functions and visualize data.
The Python code is written by the LLM using probabilistic methods. The code is then executed by a standard Python interpreter. The reliability of the final answer comes from having outsourced the computation to the verified, deterministic Python libraries. The potential unreliability comes from the LLM’s probabilistic process of writing the code that uses those libraries.
In addition, the queries may give incorrect answers if the problems are inaccurately posed. The LLM could misunderstand the prompt, have bugs in its code or fail to handle boundary cases.
Gemini and the paid version of Chat GPT give automatic user access to the Python packages NumPy, SymPy, SciPy, Pandas, Matplotlib.
6. How much do LLMs ‘know’?
Current LLMs are trained on human data, which means they cannot generate knowledge outside of that data. While they can go beyond the knowledge of any single human, they can only come up with new ideas through extrapolation, not discovery from first principles. Their core capability is Natural Language Reasoning - the ability of LLMs to process, synthesize and generate human language by identifying and replicating statistical patterns from vast amounts of data in the form of tokens. In contrast, Symbolic Deduction is manipulation based on formal rules applied to abstract symbols, as in the operation of a Computer Algebra System. Current LLMs are built to imitate reasoning based on patterns, but are incapable of performing true symbolic deduction.
However, just as mathematicians are becoming familiar with the use of LLMs, their internal architecture is already changing. In 2025, Google’s DeepMind achieved a gold-medal level score alongside the International Mathematical Olympiad using their AlphaProof and AlphaGeometry systems. These systems did not use a separate formal proof assistant. The key innovation in their new model is the integration of Symbolic Deduction with Natural Language Reasoning, representing a first step towards an AI-based alternative to formal proof assistants. Their new models also incorporate techniques like parallel thinking, which allowed them to explore multiple solutions simultaneously.
The significant advancement in reasoning capability has been achieved by an advanced form of chain-of-thought and tree-of-thought reasoning. These are important tools in when using LLMs for mathematics. We give some examples in Section 12.
The next frontier goes even further, with systems like Google’s AlphaZero, which is not based on human-generated data. AlphaZero is given only the rules and foundations of a subject. It then learns (for example, in chess) by self-play and it generates its own training data. It is therefore likely that a future ‘AlphaMath’ system could be trained only on the axioms and rules of inference of ZFC set theory. Such a system would engage in genuine ‘self-discovery’ by attempting to prove theorems. Over time, with human feedback, it would begin to recognize patterns and strategies that lead to successful proofs.
7. LLMs and patterns in higher mathematics
An LLM can identify mathematical patterns that are imperceptible to humans. It performs massive scale statistical analysis on the symbolic structure of mathematics itself. It can find cross-disciplinary correlations and potential analogies that humans may not think of. It is unconstrained by human cognitive biases and the boundaries of human knowledge.
The way mathematicians are trained to think assumes the progression of ideas of mathematical discovery. We have learned concepts in a chronological sequence. LLMs do not ‘know’ that mathematical fields developed sequentially. For an LLM, mathematical concepts coexists simultaneously. They are linked only by the statistical and structural patterns embedded in the language of mathematics. Due to their inherent design, they can propose unexpected pathways, connections and points of view.
7.1. Advantages of this capability:
An LLM sees mathematics as a language with a complex grammar and vocabulary built up of symbols and diagrams. It learns the statistical relationships between the tokens in this language.
Trained on the entire available library of mathematics and science without disciplinary boundaries, it learns the statistical relationships between tokens, allowing it to explore the entire space of mathematical connections, not just those that individual mathematicians are trained to look for.
LLMs don’t ‘see’ symbols and equations. There is a map $\varphi:V\to\mathbb{R}^{d}$ (for $d\leq 13,000$) from the vocabulary of tokens to a high-dimensional vector space. The dimension $d\sim|V|^{1/4}$ is a model parameter. Relationships in the vocabulary are translated into geometric relationships in $\mathbb{R}^{d}$.
The model can detect geometric relationships in $\mathbb{R}^{d}$ that are impossible for humans to visualize and that are not obvious from the symbolic definitions. For example, it uses a sophisticated method of projecting tokens onto subspaces via linear transformations to isolate a specific feature to look at. It then reviews all isolated facts to decide on an output.
It is this ability to perceive and analyze non-obvious geometric patterns that allows LLMs to generate novel conjectures and insights when prompted appropriately.
7.2. Human-LLM collaborative research
Using an LLM as a collaborative research partner, rather than as a search engine or text generator, involves a continuously refined, iterative dialogue. The researcher guides the model’s output through a series of increasingly specific constraints, corrections, and questions, often in a loop that involves downloading, correcting, and re-uploading the model’s work for further prompting.
For open research questions, this can be a lengthy process, sometimes involving hundreds of prompts in a single chat, to obtain a final product that is useful. A key technique is ‘in-context learning’, where researchers provide background information like papers or books via prompts. The LLM keeps this material in its short-term memory, becoming more adept at handling the research topic as the conversation history forms an expanding context window.
While not necessarily a time-saving endeavor, this collaborative process can reveal ideas and connections that lie beyond a human researcher’s own spectrum.
8. Why AI is a useful tool in combinatorial group theory
We discuss some examples from group theory. However, many of the underlying principles are applicable to other topics.
Let $G$ be a group defined by a presentation $G=\langle X\mid R\rangle$, where $X$ is a set of generators and $R$ is a set of relations. This is a compact way to define a group, but it hides immense complexity.
The word problem (determining if a word in the generators is equal to the trivial element of the group) is unsolvable in general [MT1973].
Computer algebra systems such as GAP, Magma, and Sage can solve the word problem for specific families of groups using coset rewriting methods and structural properties. For new groups whose structure is not well understood, an LLM can be used as an interactive research assistant.
A simple example: let $G=BS(1,2)=\langle x,y\mid x^{-1}yx=y^{2}\rangle$. This is the so-called Baumslag-Solitar group $BS(1,2)$. It’s word problem is known to be solvable.
A solution of the word problem involves deciding on an efficient sequence of applications of the group relations to achieve the goal of trying to get to the identity element of the group.
Solve the word problem for the word $w=yxy^{-2}x^{-1}$:
We first rewrite $x^{-1}yx=y^{2}$ as $yx=xy^{2}$. Then
${yx}y^{-2}x^{-1}={xy^{2}}y^{-2}x^{-1}=xx^{-1}=1.$
So the word $w=yxy^{-2}x^{-1}$ represents the identity element in $G$.
8.1. Solving algorithmic research questions
Algorithmic search problems in group theory, such as the word problem, conjugacy problem (deciding if two words are conjugate in a given group), and triviality problem (deciding if a group presentation presents the trivial group), are undecidable in general. Even in decidable cases, the search space for groups with many generators and intertwined relations is often astronomically large, making it intractable for hand calculation.
AI can address this by effectively simplifying words. It has strategies to navigate the search space by deciding which group relations are the most useful to apply.
However, this is not true symbolic manipulation: LLMs mimic the process through pattern recognition, so their output must always be verified and corrected.
In addition, LLMs can generate code for computer algebra systems such as GAP, Magma, and SageMath with varying degrees of reliability. Proficiency is particularly high with SageMath, given its Python-based syntax and integration with GAP for computations in group theory and other areas.
9. Examples of the use of LLMs to address pattern-matching problems in higher mathematics
You can ask an LLM to:
  • Give an explicit reference (with page numbers) on a topic known to be in some book or research paper.
  • Apply the statement of a theorem to a particular example.
  • Generate LaTex code from an uploaded pdf file of a math paper.
  • Translate a mathematics paper from another language and give the output in a LaTex file.
  • Explain a section of a physics paper in mathematical language.
  • Generate a Tikz diagram or table from a description in natural language.
  • Generate a bibliography on a specific topic.
  • Rewrite the LaTex code for an entire math paper in different notation.
  • Find the typos in an uploaded pdf file.
  • Analyze if the flow of ideas in a paper is appropriately sequential.
  • Generate LaTex, Lean, Python, Mathematica, Maple and other forms of code.
10. Comparison of LLMs
Gemini, ChatGPT and Claude all exhibit different specializations in how they activate nodes, navigate paths and narrow down possibilities within their internal graphs of statistical associations.
For the most difficult mathematical tasks, Gemini, ChatGPT, and Claude are all roughly in the same class with respect to performance.
Though Gemini models perform slightly better in advanced mathematical benchmarks, as of 2024.
Model MATH (%)
Claude 74.9
ChatGPT 76.6
Gemini 78.5
‘Advanced mathematics’ refers to the standard benchmarks for competition-level mathematics, testing symbolic reasoning in algebra, calculus, number theory, etc.
Advantages of ChatGPT are its plug-ins: WolframAlpha can be enabled in the professional version. Another plug-in called Sci-space has access to 300 million peer-reviewed journal papers.
11. General rules for writing prompts for advanced math questions:
  • Specify the task in short sentences.
  • Use ‘your task is to…’ or ‘your goal is to…’.
  • Specify the context of the task.
  • Upload all necessary background information.
  • Be explicit and detailed.
  • Include all relevant keywords.
  • Guide the reasoning process.
  • Specify the output format.
  • Verify the output.
  • Never ask for a complex proof in one shot.
  • Break a difficult task up into smaller tasks.
  • Take the model’s output and ask for it to be modified with specific constraints.
  • Constrain the method the model is allowed to use.
  • Find errors in the output and ask the model to self-correct.
You can also ask an LLM to generate a prompt for a given task.
12. Prompts for advanced mathematics
12.1. Chain-of-thought prompt
This refers to guiding the LLM to think step-by-step, generating intermediate reasoning before arriving at the final answer [IBMToT]. This type of structured prompt imposes a logical framework on the LLM’s probabilistic generation process. If the LLM shows its work, this reduces errors.
Example prompt:
Your task is to generate a detailed description of the representation of the Weyl group of $\mathfrak{sl}_{3}(\mathbb{C})$ on the dual space $\mathfrak{h}^{*}$ of its Cartan subalgebra $\mathfrak{h}$. Proceed with the following line of reasoning, evaluating each step before proceeding to the next:
[…]
12.2. Tree-of-thought prompt
This is guiding an LLM to explore multiple different deduction paths simultaneously such that it evaluates its own progress at each step and pursues the most promising path [IBMToT].
Example prompt:
Your task is to determine if the Higman group, defined by the presentation
$H=\langle a,b,c,d\mid a^{-1}ba=b^{2},\,b^{-1}cb=c^{2},\,c^{-1}dc=d^{2},\,d^{-1}ad=a^{2}\rangle$
is trivial or non-trivial. Proceed with the following line of reasoning, evaluating each step before proceeding to the next:
Is the Higman Group Trivial? Branch 1: Elementary Tests 1a. Abelianization $H_{ab}$ is trivial. No non-trivial abelian quotients. Inconclusive 1b. Direct Simplification Leads to complexity. Inconclusive Branch 2: Structural Analysis 2a. Identify group generated by individual relations Identifies non-trivial $BS(1,2)$ building blocks. 2b. Explain Construction Built via HNN extensions with embedded $BS(1,2)$ subgroups. Conclusion: The Higman Group is NON-TRIVIAL.
Figure 1. Tree-of-thought prompt diagram for non-triviality of the Higman group.
These types of example prompts have been shown to be useful in guiding an LLM how to approach complex mathematical questions.
13. Fact-checking: retrieval-augmented generation
The default output of an LLM is an output that is based on its training data. The ability to access the internet to fact-check in real time is a feature built on top of an LLM, through a mechanism called ‘retrieval-augmented generation’. It may require a specific prompt in order to invoke this feature. Users can also explicitly request a list of websites and references used.
When asked a question that requires fact-checking, Gemini automatically queries the Google Search index in real time. It then synthesizes the information from the top-ranked pages to give an answer and often provides direct links.
The current version of ChatGPT has a ‘reasoning mode’ in its new model which autonomously decides when to use its web-browsing tools. Before finalizing an answer, the model evaluates and corrects its own results. It internally grades its answer against criteria to ensure it meets a high-quality standard.
Microsoft Copilot is designed to be an AI-powered search assistant. Most queries initiate a live search on Bing and links are usually provided.
14. A single chat is a ‘context window’
An LLM’s short-term memory is its context window, which consists of a single chat session. This window contains the entire conversation history, including uploaded documents and any corrections made by the user. Its capacity is measured in tokens: while a 120K window can hold a novel, models like Gemini offer an unprecedented 1 million tokens.
However, this memory is finite. Once the token limit is reached, the LLM will begin to forget the earliest messages in the conversation. Work on a difficult mathematics problem must be contained within a single chat: the LLM needs all information and previous steps to be present simultaneously to solve a complex problem.
Since an LLM has no inherent memory between separate sessions, new chats must be initiated with any updated information and results.
15. Other ways to influence the output from an LLM
15.1. Design your outputs
In LLMs, there is a ‘system instruction’ feature which can be used to influence the type of output. You can enter a ‘pre-prompt’ with your own personal profile that governs the model’s behavior for the entire chat session.
15.2. Changing the ‘knobs’
System settings in an LLM can also be manually changed in order to make mathematical proofs more rigorous and less random.
For mathematical rigor, the most important setting is temperature’. This controls the randomness of the output by reshaping the probability distribution. A high temperature (such as 0.9) encourages more diverse and novel responses. A low temperature makes the output more deterministic. The recommended setting for math in Gemini is to lower the temperature to 0.2 or 0.1.
15.3. Advanced: system prompts
‘Back-end’ engineering, such as writing system prompts in an LLM gives more direct control and can impact the amount of detail and rigor in a proof.
16. Hybrid systems: LLMs with Computer Algebra Systems or formal proof assistants
Because LLMs are not substitutes for formal proof or verified computation, their outputs require rigorous human scrutiny. The primary way to offset this limitation is to use them in conjunction with formal tools like Computer Algebra Systems or formal proof assistants such as Lean. For example, ChatGPT can integrate with WolframAlpha. The user can explicitly prompt ChatGPT in the Wolfram chat window, or allow it to decide when to outsource a computation.
Formal proof assistants like Lean allow for the expression and mechanical verification of mathematical proofs. Lean implements a version of Dependent Type Theory known as the Calculus of Inductive Constructions. This provides an alternative to the ZFC axioms of set theory. It allows the expression of mathematical claims, mechanically checks proofs of these claims and helps to find formal proofs.
Type Theory is advantageous for formalization because it embeds mathematical meaning directly into its syntax. Its system of Types prevents false statements and builds properties directly into the definition of its objects. In ZFC set theory, many of these logical statements require separate proofs.
However, writing Lean code is notoriously difficult, requiring not just programming skill but the precise formalization of abstract mathematical ideas in Type Theory. In this framework, even small logical gaps or unstated assumptions must be made explicit and rigorously verified. Precise use of syntax is required. One of the benefits of the Lean environment is Lean’s Mathlib: a vast, open-source digital encyclopedia of formally verified mathematics.
This is where LLMs can potentially play a transformative role. Research labs are actively working to bridge this difficulty by using LLMs to write Lean code and accept inputs in natural language, making formal verification more accessible.
17. Math research alongside LLMs and Lean
This is a brief description of joint research with Ayush Khaitan and Scott H. Murray alongside LLMs and Lean [CKM]. Results have been verified using Lean 4.
The project was to study an infinite dimensional analog $S(\mathfrak{m})$ of the Steinberg group, given by generators and relations, associated to the Monster Lie algebra $\mathfrak{m}$. This is an infinite dimensional Lie algebra which is an example of a Borcherds algebra, also called a generalized Kac–Moody algebra.
We proposed a presentation for the group and used LLMs is to assist in determining properties of $S(\mathfrak{m})$.
The work was done using a combination of usual research techniques, laborious use of context prompting, chain-of-thought prompting, and iterative prompting of LLMs, particularly Gemini and ChatGPT, verification and determining errors in the output and reprompting.
17.1. New results suggested by Gemini
During a long chat with Gemini, the model suggested the following:
Working with the group presentation indicates that the group $S(\mathfrak{m})$ is a perfect group. That is, $S(\mathfrak{m})=(S(\mathfrak{m}),S(\mathfrak{m}))$ is equal to its commutator subgroup. This property holds for certain Steinberg groups of finite dimensional simple Lie algebras.
Since the setting of finite dimensional Steinberg groups is vastly different, this had not seemed like a plausible research direction before Gemini’s suggestion.
The usual method for proving that Steinberg groups of finite dimensional simple Lie algebras are perfect involves showing that each generator is contained in the commutator subgroup. The result turned out to be true for a larger group $G(\mathfrak{m})$ constructed in [CJM], containing toral elements.
Gemini also aided in completing the highly technical proof and suggested that the group $G(\mathfrak{m})$ behaves like a simple group. This is a plausible suggestion which was already being considered.
17.2. Challenging calculations
Let $L$ be a $\mathbb{Z}\oplus\mathbb{Z}$-graded free Lie algebra on a specified set of generators.
Two different LLMs were used to obtain recursive formulas for the dimensions of the homogeneous components $L_{(m,n)}$ of $L$, for $m,n>0$.
Given $(m,n)$ with $m>0,n>0$, let $P(m,n)$ be the set of all partitions of the ordered pair $(m,n)$ into a sum of the form ${\sum}_{r,s}a_{rs}(r,s)$ such that $a_{rs},r,s\in\mathbb{N}_{\geq 1}$, ${\sum}_{r,s}a_{rs}\cdot r=m$, and ${\sum}_{r,s}a_{rs}\cdot s=n$.
For example $P(3,2)=\{(3,2),(1,1)+(2,1)\}$ where $a_{3,2}(3,2)=1(3,2)$ and $a_{1,1}(1,1)+a_{2,1}(2,1)=1(1,1)+1(2,1)$.
Given the string of coefficients $(a_{r,s})$, and a partition $a\in P(m,n)$, [JurJPAA1998] and [JLW1995] gave the formula:
$d(i,j)={\sum}_{k\in{\mathbb{Z}}_{+}\atop k(m,n)=(i,j)}{1\over k}\mu(k){\sum}_{a\in P(m,n)}{(\sum a_{rs}-1)!\over\prod a_{rs}!}\prod n(r,s)^{a_{rs}}$
where $n(i,j)$ is the number of generators of the algebra $L$ of degree $(i,j)$ and $\mu(k)$ is the Möbius function.
Using a staged prompting method, Gemini and ChatGPT/WolframAlpha produced the same symbolic expressions for $d(2,j)$ in terms of $n(r,s)$, for $j$ from 1 to 50, by first computing the partitions and then the $d(i,j)$. Both LLMs also gave a recursive strategy for the calculation of $d(i,j)$ in general.