### v2

# ADVANCING MATHEMATICS RESEARCH WITH GENERATIVE AI

Abstract.
The main drawback of using generative AI models for advanced mathematics is that these models are not logical reasoning engines. However, Large Language Models, and their refinements, can pick up on patterns in higher mathematics that are difficult for humans to see. By putting the design of generative AI models to their advantage, mathematicians may use them as powerful interactive assistants that can carry out laborious tasks, generate and debug code, check examples, formulate conjectures and more. We discuss how generative AI models can be used to advance mathematics research. We also discuss their integration with Computer Algebra Systems and formal proof assistants such as Lean.
1. Introduction
Mathematicians have mixed views about the role of generative AI models in the mathematical landscape. While these models, such as Large Language Models (LLMs), can look convincingly like they replicate known mathematics, on careful scrutiny, it is apparent that they are just masters of the rhetoric of mathematics. They don’t meet the standards of rigor that the field requires.
This shortfall is built into the design of LLMs. They are fundamentally statistical, not logical, engines. As next-word predictors, they lack a built-in engine for formal symbolic deduction, meaning their emergent ‘reasoning’, called Natural Language Reasoning, is derived from statistical patterns in their training data.
However, this statistical foundation also provides certain benefits, as LLMs can pick up on patterns in higher mathematics that are difficult for humans to see. In particular, they have a learned geometric representation of mathematical language. By leveraging these capabilities and working with LLMs as they were designed to operate, mathematicians can use them as powerful interactive assistants.
The AI landscape is rapidly evolving beyond standard LLMs, with Large Reasoning Models (LRMs) and Large Context Models (LCMs) emerging as the next generation of this technology. While built on the foundation of LLMs, these models have distinct architectures and methods of training (see Section 4).
In Section 2, we discuss the mechanics of probabilistic language generation. In Section 3, we discuss how to manage the context window. In Sections  59, we describe the capabilities and limitations of LLMs, the nature of their training data, and their ability to detect geometric patterns in high-dimensional representations of mathematical language. In Sections 1016, we discuss applications and user techniques. This includes examples from combinatorial group theory, a comparison of current models, a guide to prompt engineering, and fact-checking. In Section 17 we discuss the integration of LLMs with Computer Algebra Systems (CAS) and formal proof assistants like Lean.
2. The mechanics of probabilistic language generation
2.1. The LLM as a probabilistic knowledge graph
Consider the following simplified view of a generative AI model as a ‘knowledge graph’ which represents all possible connections and probabilities within the LLM’s knowledge base.
Each word is a node in the graph. Each edge is a statistical relationship or an association between words, which cluster to form concepts.
additionnegative numbersirrational numbersintegersreal numbersinverseCONVERGED PROBABILITY
Weighted edges between trillions of nodes represent how words or concepts are related. The weights are probabilities indicating the likelihood of connection.
An LLM generates text by predicting the next word, starting from the prompt, while calculating the probabilities of all possible next nodes. It then chooses one of the most likely paths through the knowledge graph.
2.2. Tokens: the building blocks of language
In AI, a token is the fundamental building block of language. AI models convert input materials (prompts) into tokens. A token $t_{i}$ is a digital representation of a word, a symbol, punctuation, a diagram, or other input.
Let $V$ be a finite vocabulary consisting of token representations of the training data for an LLM. An input to an LLM is a sequence of tokens $(t_{1},t_{2},\dots,t_{n})$. An output is a sequence of tokens $(t_{1},t_{2},\dots,t_{m})$, where each $t_{i}\in V$.
The fundamental operation of an LLM is to compute the probability of the next token, $t_{k+1}$, conditioned on the sequence of all preceding tokens $(t_{1},t_{2},\dots,t_{k})$.
All AI language models have a maximum token limit for input and output. The token input limit has an impact on the work possible in a single chat session.
2.3. How the prompt shapes an LLM’s output
The prompt provides an initial sequence of tokens that the model uses to condition its output. The LLM generates text by sequentially examining the existing sequence of tokens, assigning a probability to every possible next token, and then choosing one to add to the sequence.
The probability of each potential next token is calculated based on the full sequence of tokens that came before it – both the original prompt and any text already generated.
The final output is the result of a chain of these probabilistic decisions. The framing of the initial prompt is critical in determining the usefulness of the output.
The design of prompts, or ‘prompt engineering’, significantly influences the behavior of an LLM [PromptGuide], [PromptEngineering]. It disambiguates the input, leading to more relevant and useful output.
3. A single chat is a ‘context window’
A generative AI model’s short-term memory is its context window, which consists of a single chat session. This window contains the entire conversation history, including uploaded documents and any corrections made by the user. Its capacity is measured in tokens: while a 120K window can hold a novel, models like Gemini 2.5 Pro offer an unprecedented 1 million tokens.
However, this memory is finite. Once the token limit is reached, the AI model will begin to forget the earliest messages in the conversation. Work on a difficult mathematics problem must be contained within a single chat: the AI model needs all information and previous steps to be present simultaneously to solve a complex problem.
Since a generative AI model has no inherent memory between separate sessions, new chats must be initiated with any updated information and results.
4. Large Reasoning Models and Large Context Models
Large Reasoning Models (LRMs), also known as a Reasoning Language Model (RLMs), are a class of LLMs trained to carry out tasks that require multi-step deduction, critical thinking, and structured problem-solving. LRMs ‘think’ before they respond.
They deliberate, explore various solution paths, evaluate intermediate steps and revise their reasoning.
LRMs integrate frameworks that support reasoning structures, such as ‘chain-of-thought’ and ‘tree-of-thought’ structures into their systems. We give some examples in Section 14.
The training of LRMs diverges from that of LLMs. They utilize a technique called ‘process supervision’. Unlike the training process of LLMs that only rewards a correct final answer, process supervision rewards the correctness of the intermediate reasoning steps.
Large Context Models (LCMs) are defined by their large context windows and the volume of information that they can handle. This allows them to accept and analyze large datasets, entire codebases, or multimodal inputs, without requiring external retrieval systems. Their training data includes long-form content to improve their ability to track long-range dependencies between input tokens.
These advanced capabilities are becoming integrated into leading AI platforms. The cost, however, is additional computing power and increased time for responses.
Google’s Gemini models, particularly the 2.5 series, demonstrate both LRM and LCM characteristics. The models also use internal ‘thinking processes’ for improved reasoning.
ChatGPT has a ‘reasoning mode’ in its model OpenAI o3. Reinforcement learning was used to teach OpenAI o3 to ‘think’ before generating answers, using what OpenAI refers to as ‘private chain-of-thought’ processes built into their system prompts. This allows the model to plan ahead and reason through tasks, performing a series of intermediate reasoning steps to assist in solving the problem.
Claude models use ‘extended thinking’ or the ability to ‘think out loud’. As LCMs, they offer substantial context windows
5. Drawbacks of using generative AI in mathematics
The primary drawback of using current AI models for mathematics is that they are probabilistic pattern-matchers, not logical reasoning engines. Their core function is to predict the next most likely word or symbol in a sequence, not to apply deterministic mathematical rules. Generative AI models are experts at reconstructing what a proof should ‘look like’. This limitation leads to several key problems:
Mathematical hallucinations: AI models generate outputs that look plausible and are often formatted correctly but contain nonsensical logic, invented theorems, or critical errors. They can state these falsehoods with the same confident authority as factual information.
Lack of true reasoning: Unlike a Computer Algebra System, an LLM cannot manipulate symbolic expressions according to deterministic rules. This makes it prone to subtle but serious errors in algebra, analysis, arithmetic, and logic. It is fundamentally incapable of handling verification.
Propagation of errors: AI models learn from their training data, which includes errors and misconceptions found online. For example, it is known that ChatGPT 4.0 contains training data from retracted scientific papers [MITTR2025]. An LLM will inherit and reproduce these mistakes, making them unreliable for tasks that require rigor.
Consequently, mathematical output from an LLM has to be carefully scrutinized for logical and symbolic correctness.
6. Mathematical training data for generative AI models
Training data for generative AI models includes web documents, code with mathematical content, textbooks, online faculty-authored lecture notes and course materials, solutions to problem sets, as well as academic and scientific journal papers and content from arXiv. Google’s data collection from Google Books, for example, is unparalleled in its scale.
The training data includes the standard undergraduate mathematics curriculum from US and other universities, as well as the standard coursework curriculum for PhD programs in mathematics from US and other universities.
Differences in the content and handling of training data of AI companies are primary reasons why various AI models have distinct strengths, weaknesses, and ‘personalities’, especially in a specialized domain like mathematics. Curation and filtering of data, such as removing low-quality content and emphasizing trusted sources, is essential. However, the training material for each model and its handling remains a closely guarded trade secret.
A key limitation is that a generative AI model’s knowledge base is heavily biased towards materials that are easily digitized.
7. Generative AI models simulate computer algebra systems
Advanced generative AI models simulate a scientific computing environment by delegating the computation to specialized Python libraries.
Table 1. Key Python libraries used by generative AI models for mathematics
Library Main purpose Key capabilities & LLM use
NumPy Numerical Computing Handles arrays & matrices. Used by generative AI models for linear algebra (matrix multiplication, row reduction, eigenvalues).
SymPy Symbolic Mathematics Performs precise algebra & calculus. Used for symbolic derivatives, integrals, and simplifying expressions.
SciPy Scientific Computing Provides advanced numerical routines. Used for optimization (finding minima/maxima) and solving differential equations.
Pandas Data Analysis Manages structured, table-like data. Used for reading and analyzing data from files such as Excel.
Matplotlib Plotting & Visualization Creates a wide variety of 2D graphs and charts. Used to plot functions and visualize data.
The Python code is written by the LLM using probabilistic methods. The code is then executed by a standard Python interpreter. The reliability of the final answer comes from having outsourced the computation to the verified, deterministic Python libraries. The potential unreliability comes from the LLM’s probabilistic process of writing the code that uses those libraries.
In addition, the queries may give incorrect answers if the problems are inaccurately posed. The LLM could misunderstand the prompt, have bugs in its code or fail to handle boundary cases.
Gemini and the paid version of ChatGPT give automatic user access to the Python packages NumPy, SymPy, SciPy, Pandas, Matplotlib.
8. How much do generative AI models ‘know’?
Current generative AI models are trained on human data, which means they cannot generate knowledge outside of that data. While they can go beyond the knowledge of any single human, they can only come up with new ideas through extrapolation, not discovery from first principles. Their core capability is Natural Language Reasoning - the ability of generative AI models to process, synthesize and generate human language by identifying and replicating statistical patterns from vast amounts of data in the form of tokens.
In contrast, Symbolic Deduction is manipulation based on formal rules applied to abstract symbols, as in the operation of a Computer Algebra System. Current generative AI models are built to imitate reasoning based on patterns, but are incapable of performing true symbolic deduction.
However, just as mathematicians are becoming familiar with the use of generative AI models, their internal architecture is already changing.
In 2024, Google’s DeepMind reported silver-medal level on International Mathematical Olympiad problems using AlphaProof and AlphaGeometry. In 2025, DeepMind reported gold-medal level using Gemini ‘Deep Think’. These systems did not use a separate formal proof assistant. The key innovation in their new model is the integration of Symbolic Deduction with Natural Language Reasoning, representing a first step towards an AI-based alternative to formal proof assistants. Their new models also incorporate techniques like parallel thinking, which allowed them to explore multiple solutions simultaneously.
The significant advancement in reasoning capability has been achieved by an advanced form of chain-of-thought and tree-of-thought reasoning. These are important tools when using generative AI models for mathematics. We give some examples in Section 14.
The next frontier goes even further, with systems like Google’s AlphaZero, which is not based on human-generated data. AlphaZero is given only the rules and foundations of a subject. It then learns (for example, in chess) by self-play and it generates its own training data. It is therefore likely that a future ‘AlphaMath’ system could be trained only on the axioms and rules of inference of ZFC set theory. Such a system would engage in genuine ‘self-discovery’ by attempting to prove theorems. Over time, with human feedback, it would begin to recognize patterns and strategies that lead to successful proofs.
9. generative AI models and patterns in higher mathematics
A generative AI model can identify mathematical patterns that are imperceptible to humans. It performs massive scale statistical analysis on the symbolic structure of mathematics itself. This allows it to simulate a mathematician’s intuition. It can find cross-disciplinary correlations and potential analogies that humans may not think of. It is unconstrained by human cognitive biases and the boundaries of human knowledge.
The way mathematicians are trained to think assumes the progression of ideas of mathematical discovery. We have learned concepts in a chronological sequence. AI models do not ‘know’ that mathematical fields developed sequentially. For an AI model, mathematical concepts coexist simultaneously. They are linked only by the statistical and structural patterns embedded in the language of mathematics. Due to their inherent design, they can propose unexpected pathways, connections and points of view.
9.1. Advantages of this capability:
A generative AI model sees mathematics as a language with a complex grammar and vocabulary composed of symbols and diagrams. It learns the statistical relationships between the tokens in this language.
Trained on the entire available library of mathematics and science without disciplinary boundaries, an LLM learns the statistical relationships between tokens, allowing it to explore the entire space of mathematical connections, not just those that individual mathematicians are trained to look for.
Generative AI models don’t ‘see’ symbols and equations. There is a map $\varphi:V\to\mathbb{R}^{d}$ (for $d\leq 13,000$) from the vocabulary $V$ of tokens to a high-dimensional vector space (the embedding space). The dimension $d\sim|V|^{1/4}$ is a model parameter. Relationships in the vocabulary are translated into geometric relationships in $\mathbb{R}^{d}$ [mikolov2013].
Vocabulary VVmanwomanqueenkingEmbedding Space $\mathbb{R}^{d}$$\varphi(\text{man})$$\varphi(\text{king})$$\varphi(\text{woman})$$\varphi(\text{queen})$$\vec{v}_{\text{royal}}$$\vec{v}_{\text{royal}}$$\vec{v}_{\text{gender}}$$\vec{v}_{\text{gender}}$$\varphi:V\to\mathbb{R}^{d}$
The model can detect geometric relationships in $\mathbb{R}^{d}$ that are impossible for humans to visualize and that are not obvious from the symbolic definitions.
The image $\varphi(w)$ of a word $w$ doesn’t have a single location in the embedding space. Its ‘meaning’ is its relationship to all other images of tokens in the space and their locations. Furthermore, this meaning changes with context. The model applies a linear transformation $T$ to $\varphi(w)\in\mathbb{R}^{d}$ to project it onto a subspace relevant to its contextual role. This allows it to isolate a specific feature to look at. It then reviews all isolated facts and uses piecewise linear maps to put these facts together and compile an output.
This ability to perceive and analyze non-obvious geometric patterns allows the AI model to generate novel conjectures and insights when suitably prompted.
9.2. Human-LLM collaborative research
Using a generative AI model as a collaborative research partner, rather than as a search engine or text generator, involves an iterative dialogue that must be repeatedly corrected and refined. The user can improve the model’s output through a series of increasingly specific constraints, corrections, and questions, often in a loop that involves downloading, correcting, and re-uploading the model’s work for further prompting.
For open research questions, this can be a lengthy process, sometimes involving hundreds of prompts in a single chat, to obtain a final product that is useful. A key technique is ‘in-context learning’, where researchers provide background information like papers or books via prompts. The LLM keeps this material in its short-term memory. It becomes more adept at handling the research topic as the conversation history forms an expanding context window.
While not necessarily a time-saving endeavor, this collaborative process can reveal ideas and connections that lie beyond a human researcher’s own spectrum.
10. Why AI is a useful tool in combinatorial group theory
We discuss some examples from group theory. However, many of the underlying principles are applicable to other topics.
Let $G$ be a group defined by a presentation $G=\langle X\mid R\rangle$, where $X$ is a set of generators and $R$ is a set of relations. This is a compact way to define a group, but it hides immense complexity.
The word problem (determining if a word in the generators is equal to the trivial element of the group) is unsolvable in general [MT1973].
Computer Algebra Systems such as GAP, Magma, and SageMath can solve the word problem for specific families of groups using rewriting methods and structural properties.
A simple example: let $G=BS(1,2)=\langle x,y\mid x^{-1}yx=y^{2}\rangle$. This is the so-called Baumslag-Solitar group $BS(1,2)$. Its word problem is known to be solvable.
A solution of the word problem involves deciding on an efficient sequence of applications of the group relations to achieve the goal of trying to get to the identity element of the group.
As an example, we solve the word problem for the word $w=yxy^{-2}x^{-1}$:
We first rewrite $x^{-1}yx=y^{2}$ as $yx=xy^{2}$. Then
${yx}y^{-2}x^{-1}={xy^{2}}y^{-2}x^{-1}=xx^{-1}=1.$
Thus the word $w=yxy^{-2}x^{-1}$ represents the identity element in $G$.
10.1. Solving algorithmic research questions
Algorithmic search problems in group theory, such as the word problem, conjugacy problem (deciding if two words are conjugate in a given group), and triviality problem (deciding if a group presentation presents the trivial group), are undecidable in general. Even in decidable cases, the search space for groups with many generators and intertwined relations is often astronomically large, making it intractable for hand calculation.
AI can address this by effectively simplifying words. It has strategies to navigate the search space by deciding which group relations are the most useful to apply.
However, this is not true symbolic manipulation: generative AI models mimic the process through pattern recognition, so their output must always be verified and corrected.
In addition, generative AI models can generate code for Computer Algebra Systems such as GAP, Magma, and SageMath with varying degrees of reliability. Proficiency is particularly high with SageMath, given its Python-based syntax and integration with GAP for computations in group theory and other areas.
11. Examples of the use of generative AI models to address pattern-matching problems in higher mathematics
You can ask a generative AI model to:
  • Give an explicit reference (with page numbers) on a topic known to be in some book or research paper.
  • Apply the statement of a theorem to a particular example.
  • Generate LaTeX code from an uploaded pdf file of a math paper.
  • Translate a mathematics paper from another language and give the output in a LaTeX file.
  • Explain a section of a physics paper in mathematical language.
  • Generate a Tikz diagram or table from a description in natural language.
  • Generate a bibliography on a specific topic.
  • Rewrite the LaTeX code for an entire math paper in different notation.
  • Find the typos in an uploaded pdf file.
  • Analyze if the flow of ideas in a paper is appropriately sequential.
  • Generate LaTex, Lean, Python, Mathematica, Maple and other forms of code.
12. Comparison of generative AI models
Gemini, ChatGPT and Claude all exhibit different specializations in how they activate nodes, navigate paths and narrow down possibilities within their internal graphs of statistical associations.
For the most difficult mathematical tasks, Gemini, ChatGPT, and Claude are all roughly in the same class with respect to performance. Though Gemini models performed slightly better in the standard benchmarks for competition-level mathematics, testing symbolic reasoning in algebra, calculus and number theory in 2024.
Model MATH (%)
Claude 74.9
ChatGPT 76.6
Gemini 78.5
ChatGPT integrates features like WolframAlpha which can be enabled in the professional version. Another GPT called SciSpace has access to hundreds of millions of peer-reviewed journal papers.
13. General rules for writing prompts for advanced math questions:
  • Specify the task in short sentences.
  • Use ‘your task is to…’ or ‘your goal is to…’.
  • Specify the context of the task.
  • Upload all necessary background information.
  • Be explicit and detailed.
  • Include all relevant keywords.
  • Guide the reasoning process.
  • Specify the output format.
  • Verify the output.
  • Never ask for a complex proof in one shot.
  • Break a difficult task up into smaller tasks.
  • Take the model’s output and ask for it to be modified with specific constraints.
  • Constrain the method the model is allowed to use.
  • Find errors in the output and ask the model to self-correct.
You can also ask a generative AI model to generate a prompt for a given task.
14. Structured prompts and system instructions
Structured prompts impose a logical framework on an LLM’s probabilistic generation process. The following examples of prompts have been shown to be useful in guiding generative AI models on how to approach complex mathematical questions [kojima2022]. These methods are also used internally in advanced LRM models like Gemini and ChatGPT.
14.1. Chain-of-thought reasoning
This refers to guiding the LLM to think step-by-step, generating intermediate reasoning before arriving at the final answer [IBMToT]. If the LLM shows its work, this reduces errors.
Example prompt:
Your task is to generate a detailed description of the representation of the Weyl group of $\mathfrak{sl}_{3}(\mathbb{C})$ on the dual space $\mathfrak{h}^{*}$ of its Cartan subalgebra $\mathfrak{h}$. Proceed with the following line of reasoning, evaluating each step before proceeding to the next:
[…]
14.2. Tree-of-thought prompt
This refers to guiding a generative AI model to explore multiple different deduction paths simultaneously such that it evaluates its own progress at each step and pursues the most promising path [IBMToT].
Example prompt:
Your task is to determine if the Higman group, defined by the presentation
$H=\langle a,b,c,d\mid a^{-1}ba=b^{2},\,b^{-1}cb=c^{2},\,c^{-1}dc=d^{2},\,d^{-1}ad=a^{2}\rangle$
is trivial or non-trivial. Proceed with the following line of reasoning, evaluating each step before proceeding to the next:
Is the Higman Group Trivial? Branch 1: Elementary Tests 1a. Abelianization $H_{ab}$ is trivial. No non-trivial abelian quotients. Inconclusive 1b. Direct Simplification Leads to complexity. Inconclusive Branch 2: Structural Analysis 2a. Identify group generated by individual relations Identifies non-trivial $BS(1,2)$ building blocks. 2b. Explain Construction Built via HNN extensions with embedded $BS(1,2)$ subgroups. Conclusion: The Higman Group is NON-TRIVIAL.
Figure 1. Tree-of-thought prompt diagram for non-triviality of the Higman group.
The emergence of LRMs and LCMs (Section 4) has altered the practice of prompt engineering.
For LRMs, the need for prompt engineering techniques designed to force reasoning, such as explicit chain-of-thought prompting, is reduced. The reasoning process is handled internally. Simpler, goal-oriented prompts that clearly define the problem may be more effective.
With LCMs, the focus is shifting to ‘context engineering’. The challenge becomes optimizing the content and structure of the information provided in the prompt. This maximizes performance and prevents recall accuracy from decreasing as the context window fills.
15. Fact-checking: retrieval-augmented generation
The default output of a generative AI model is an output that is based on its training data. The ability to access the internet to fact-check in real time is a feature built on top of an LLM, through a mechanism called ‘retrieval-augmented generation’, or RAG. It may require a specific prompt in order to invoke this feature. Users can also explicitly request a list of websites and references used.
When asked a question that requires fact-checking, Gemini automatically queries the Google Search index in real time. It then synthesizes the information from the top-ranked pages to give an answer and often provides direct links.
The current version of ChatGPT has a ‘reasoning mode’ in its new model which autonomously decides when to use its web-browsing tools. Before finalizing an answer, the model evaluates and corrects its own results. It internally grades its answer against criteria to ensure it meets a high-quality standard.
Microsoft Copilot is designed to be an AI-powered search assistant. Most queries initiate a live search on Bing and links are usually provided.
16. Other ways to influence the output from a generative AI model
16.1. Design your outputs
In generative AI models, there is a ‘system instruction’ feature which can be used to influence the type of output. You can enter a ‘pre-prompt’ with your own personal profile that governs the model’s behavior for the entire chat session.
16.2. Changing the ‘knobs’
System settings in certain generative AI models can also be manually changed in order to make mathematical proofs more rigorous and less random.
For mathematical rigor, the most important setting is ‘temperature’. This controls the randomness of the output by reshaping the probability distribution. For research questions and formulating conjectures, a high temperature (such as 0.9) is preferable, as it encourages more diverse and novel responses. A low temperature (such as 0.2 or 0.1) makes the output more deterministic.
Table 2. Optimized settings in Google AI studio for research in mathematics
Temperature Medium-High ($0.7$ - $1.0$) A higher temperature encourages exploration.
Top P High ($0.95$ - $1.0$) Allows the model to consider a wider, more diverse set of ‘next words’.
Thinking mode Advanced Research problems require the deepest level of reasoning.
Set thinking budget Maximum Gives the model time to explore the ‘search space’.
Code execution On Crucial for running checks with Python tools such as SymPy.
Grounding (Google Search) On Allows the model to search outside its training data.
URL Context On Allows you to give the model specific URLs as context prompts.
Structured output Off Allows unstructured brainstorming.
Function calling On Connect to external research tools such as a university library API or a computational algebra system like SageMath or WolframAlpha.
Media resolution High Essential for correct reading of symbols in handwritten notes or complicated diagrams.
Output Length Maximum Allows for an unencumbered output.
16.3. Advanced: system prompts
‘Back-end’ engineering, such as writing system prompts in a generative AI model gives more direct control and can impact the amount of detail and rigor in a proof.
17. Hybrid systems: generative AI models with Computer Algebra Systems or formal proof assistants
Because generative AI models are not substitutes for formal proof or verified computation, their outputs require rigorous human scrutiny. The primary way to offset this limitation is to use them in conjunction with formal tools like Computer Algebra Systems or formal proof assistants such as Lean. For example, the professional version of ChatGPT can integrate with WolframAlpha. The user can explicitly prompt ChatGPT in the Wolfram chat window, or allow it to decide when to outsource a computation.
It is an active area of research to develop tools that link generative AI models with Computer Algebra Systems. One such tool is Google DeepMind’s AlphaGeometry. See also [KG2025].
Formal proof assistants like Lean allow for the expression and mechanical verification of mathematical proofs. Lean implements a version of Dependent Type Theory known as the Calculus of Inductive Constructions. This provides an alternative to the ZFC axioms of set theory. It allows the expression of mathematical claims, mechanically checks proofs of these claims and helps to find formal proofs.
Type Theory is advantageous for formalization because it embeds mathematical meaning directly into its syntax. Its system of Types prevents false statements and builds properties directly into the definition of its objects. In ZFC set theory, many of these logical statements require separate proofs.
However, writing Lean code is notoriously difficult, requiring not just programming skill but the precise formalization of abstract mathematical ideas in Type Theory. In this framework, even small logical gaps or unstated assumptions must be made explicit and rigorously verified. Precise use of syntax is required. One of the benefits of the Lean environment is Mathlib: a vast, open-source digital encyclopedia of formally verified mathematics.
Research labs are actively working on ways to make the task of generating Lean code easier: from using generative AI models to write Lean code, to developing interfaces that accept inputs in natural language.
18. Should I collaborate with a generative AI model?
As we have discussed, generative AI, on its own, should not be viewed as an oracle or an authoritative source for mathematical output. Verification of output is a necessary part of using an AI model for mathematics. But it can be used as a creative collaborator that provides ideas, perspective and potential new directions. When combined with a Computer Algebra System, it can reliably perform laborious tasks and calculations. When paired with formal verification tools, it transforms the research process itself.
The author of this work now has several LLM+CAS+Lean-assisted research projects in infinite dimensional algebra and group theory. The use of generative AI models has been particularly revealing. There have been several instances where AI models have suggested novel research directions that had initially been dismissed as implausible or irrelevant. Yet, these suggestions turned out to yield fruitful and unexpected paths forward.